# date21082024

### v1

@@ Abstract
Introduction.—Resonance is a fundamental and ubiquitous physical phenomenon that plays a key role in fields as diverse as quantum mechanics [1], structural engineering [2], electromagnetics [3], and fluid dynamics [4]. In closed systems, resonances are associated with normal modes, which correspond to the orthogonal eigenvectors of the system’s Hermitian Hamiltonian. The associated real eigenvalues define the spectrum of permissible oscillation frequencies or energy levels [5]. Solutions of eigenvalue problems, however, are not in general analytically possible. Approximate methods, in which the system of interest is modelled by perturbing a solvable reference problem, are therefore frequently employed. Rayleigh-Schrödinger perturbation theory, in which the perturbed solution is written as a series expansion over all unperturbed eigenstates, is perhaps the most famous approach [6], however a multitude of alternate theories have been developed. Brillouin-Wigner perturbation theory, for example, can provide better convergence properties for larger perturbations if the spectrum comprises of mostly bound states [7]. Alternatively, many-body perturbation theory focuses on large numbers of particles, whereby particle interactions are considered as a perturbation to a non-interacting system [8].
@@
Although closed systems are useful theoretical ideals, in reality a system interacts with its external environment. The coupling of normal modes to external degrees of freedom introduces potential loss channels, which can alter the nature and energy of resonant modes. So-called open, or scattering, systems can no longer be described by a Hermitian Hamiltonian and a coupling operator must be introduced to form an effective non-Hermitian Hamiltonian [9]. Alternatively, one can use the closely related scattering matrix, $\mathbf{S}$, which derives from the resolvent of the effective Hamiltonian [10]. The resonant states of the system, now termed quasi-normal modes, are then identified from the poles of the scattering matrix when analytically continued into the complex frequency plane [11, 5]. Loss to the environment pushes the resonant frequencies off the real axis, resulting in complex-valued eigenfrequencies $\omega_{p}$, where $(\omega_{p})$ and $-\mathrm{Im}(\omega_{p})/2$ are the resonant frequency and linewidth respectively.
@@
Open systems are known to possess a number of interesting resonant phenomena and scattering anomalies, such as coherent perfect absorption [12], exceptional points [13], and bound states in the continuum [14]. Moreover, resonances in open systems are susceptible to environmental variations, such as changes in temperature, external electromagnetic fields or material composition, a concept underpinning many sensing technologies [15, 16, 17]. Perturbed non-Hermitian scattering systems can exhibit additional unique behavior, such as anomalous resonance shifts [18, 19] and nonlinear sensitivity at exceptional point resonances [20]. Evaluation of resonant modes and their properties therefore remains an important task in, for example, evaluating state lifetimes, laser dynamics and sensor sensitivity. Traditional perturbation theories, however, are not directly applicable to open systems due to mathematical complexities introduced by quasi-normal modes, which lack orthogonality and grow exponentially away from system [21]. Resonances with high quality factors only couple weakly to the environment, such that these effects are minor. In the electromagnetic domain, it has been shown that in such cases resonance shifts can be calculated in terms of changes in the stored electromagnetic energy [22, 23]. The validity of such expressions however is limited, and, in the presence of significant losses, more careful consideration of the normalization of quasi-normal modes [24] or alternative mode expansions [25] become necessary. In this paper, we present a novel formulation for determining resonance shifts due to arbitrary system perturbations based on generalized Wigner-Smith operators. In the case of high quality factor resonances, we show that these operators are related to a system’s stored and dissipated energy, in full agreement with results from traditional perturbation theory. Our theory is tested against numerical simulations of resonant scattering in random photonic networks. In an accompanying paper [26], we apply our theory to low quality factor resonances and present further numerical examples.
@@
Perturbation theory.—An important tool commonly employed in the analysis of scattering systems is the Wigner-Smith time delay matrix
@@
where $\partial_{\omega}$ denotes the partial derivative with respect to frequency $\omega$ [27, 28]. $\mathbf{Q}_{\omega}$ has several interesting properties relevant to resonant scattering. For unitary systems at real $\omega$, $\mathbf{Q}_{\omega}$ is Hermitian and its eigenvalues are associated with well-defined time delays experienced by narrow-band, transient system excitations. Evaluated at scattering matrix poles, these time delays coincide with the lifetimes of the associated resonances, demonstrating a link between $\mathbf{Q}_{\omega}$ and a system’s quasi-normal modes [29]. In electromagnetic theory, it has also been shown that diagonal elements of $\mathbf{Q}_{\omega}$ can be expressed as energy-like integrals over the extent of the system [30], highlighting the connection to mode volume. By itself, however, $\mathbf{Q}_{\omega}$ lacks the specificity required to efficiently capture localized or parametric system perturbations. For this reason, we also consider the so-called generalized Wigner-Smith operator $\mathbf{Q}_{\alpha}$, defined by Eq. (1) with the replacement $\omega\rightarrow\alpha$, where $\alpha$ is an arbitrary variable. Operators of this kind have recently attracted significant attention in the optical domain as a tool for engineering light in complex scattering environments [31, 32, 33]. Importantly, by virtue of its generalized nature, $\mathbf{Q}_{\alpha}$ allows for more targeted descriptions of specific features or components of a system. For example, if $\alpha$ represents the refractive index of an isolated scatterer, $\mathbf{Q}_{\alpha}$ can be expressed in terms of the integrated optical intensity within the scatterer volume [32].
The generalized Wigner-Smith operator contains sufficient information to predict resonance shifts induced by arbitrary system perturbations. To show this, we consider a generic scattering system described by a scattering matrix $\mathbf{S}$, whose poles are assumed to be of order one. Let $\alpha$ denote some system parameter, such as the temperature, shape or size of some component of the system. If $\omega_{p}$ denotes a complex resonant frequency and the system is perturbed so that $\alpha$ changes value to $\alpha^{\prime}$, the resonant frequency will shift to $\omega_{p}^{\prime}$. To determine the change in the resonant frequency $\Delta\omega_{p}=\omega_{p}^{\prime}-\omega_{p}$ that is induced by the perturbation $\Delta\alpha=\alpha^{\prime}-\alpha$, consider the function $f(\omega,\alpha)=\det[\mathbf{S}^{-1}(\omega,\alpha)]$, whose zeros coincide with the poles of $\mathbf{S}$. Assuming the perturbation is small, a first order, multivariate Taylor expansion of $f$ gives
@@
where tr denotes the trace operator and $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}$ denotes the residue at the pole $\omega_{p}$. As can be seen from Eq. (3), for a given perturbation $\Delta\alpha$, the pole shift $\Delta\omega_{p}$ is completely determined by $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\alpha})$, which, we emphasize, should be evaluated for the unperturbed network. We also observe that the form of Eq. (3) implies the ratio of the residues of the Wigner-Smith operators effectively behaves as a condition number for $\omega_{p}$ when thought of as a function of $\alpha$ [35] . Eq. (3) is our key result, which, importantly, was derived on purely mathematical grounds, and is therefore applicable to a broad range of physical scenarios.
@@
To gain physical insight into Eq. (3), we consider now a class of optical systems composed of arbitrary, finite interior regions coupled to the environment by a finite number of dielectric waveguides. Examples of such systems include fiber Bragg resonators [36], complex networks [37] and ring resonators [38]. Note that Eq. (3) can be written in the limit form
@@
$\Delta\omega_{p}=\lim_{\omega\to\omega_{p}}\Bigg{(}-\Delta\alpha\frac{\mbox{tr}(\mathbf{Q}_{\alpha})}{\mbox{tr}(\mathbf{Q}_{\omega})}\Bigg{)},$ (4)
@@
and so the pole shift is given approximately by the expression within the limit evaluated at $\omega$ close to $\omega_{p}$. We assume for simplicity that the system permittivity $\epsilon$ is real and isotropic and that the permeability $\mu_{0}$ is that of free space. Assuming also that the resonances under consideration have high quality factors, such that $\mathrm{Im}(\omega)\ll(\omega)$, it is possible to show that [34]
$\mathbf{Q}_{\xi}=\bigg{(}\mathbf{I}+2\mathrm{Im}(\omega)\int_{\Omega}(\epsilon\mathbf{U}^{e}+\mu_{0}\mathbf{U}^{m})\,\mathrm{d}V\bigg{)}^{-1}\bigg{(}-\int_{\Omega}[\partial_{\xi}(\omega\epsilon)\mathbf{U}^{e}+\partial_{\xi}\omega\mu_{0}\mathbf{U}^{m}+2i\mathrm{Im}(\omega)(\epsilon\mathbf{V}_{\xi}^{e}+\mu_{0}\mathbf{V}_{\xi}^{m})]\,\mathrm{d}V\bigg{)},$ (5)
@@
where $\xi$ represents either $\omega$ or $\alpha$ and $\Omega$ denotes the volume occupied by the system with a boundary $\partial\Omega$ perforated only by the coupling waveguides. In Eq. (5), $\mathbf{I}$ is the identity matrix and $\mathbf{U}^{e},\mathbf{U}^{m},\mathbf{V}^{e}_{\xi}$, and $\mathbf{V}^{m}_{\xi}$ are matrices whose $(q,p)$–th elements are given by $U^{e}_{qp}=\frac{1}{4}\mathbf{E}_{p}\cdot\mathbf{E}^{*}_{q}$, $U^{m}_{qp}=\frac{1}{4}\mathbf{H}_{p}\cdot\mathbf{H}^{*}_{q}$, $V^{e}_{\xi,qp}=\frac{1}{4}\partial_{\xi}\mathbf{E}_{p}\cdot\mathbf{E}^{*}_{q}$, and $V^{m}_{\xi,qp}=\frac{1}{4}\partial_{\xi}\mathbf{H}_{p}\cdot\mathbf{H}^{*}_{q}$ respectively, where the fields $\mathbf{E}_{p}$ and $\mathbf{H}_{p}$ ($\mathbf{E}_{q}$ and $\mathbf{H}_{q}$) are those that exist throughout $\Omega$ when the system is illuminated by the $p$–th ($q$–th) incident field. Here, $p$ and $q$ should be understood as enumerating all modes in all waveguides that connect the system to the environment.
@@
As discussed in the Supplemental Material [34], the form of $\mathbf{Q}_{\xi}$ in Eq. (5) is linked to the factorization $\mathbf{Q}_{\xi}=\mathbf{A}^{-1}\mathbf{B}_{\xi}$, where $\mathbf{A}=\mathbf{S}^{\dagger}\mathbf{S}$ and $\mathbf{B}_{\xi}=-i\mathbf{S}^{\dagger}\partial_{\xi}\mathbf{S}$. We shall analyze these factors separately. First, to further simplify matters, we consider the case where the system has a single coupling waveguide supporting a single mode, so that the scattering matrix reduces to an effective scalar reflection coefficient $r_{\mathrm{eff}}$. Correspondingly, the Wigner-Smith operator $\mathbf{Q}_{\xi}$ reduces to the scalar quantity $Q_{\xi}=A^{-1}B_{\xi}$, where $A=|r_{\mathrm{eff}}|^{2}$ and $B_{\xi}=-ir^{*}_{\mathrm{eff}}\partial_{\xi}r_{\mathrm{eff}}$. To evaluate these expressions at the complex frequency $\omega$, we employ a useful transformation to recast the problem into one at a real frequency, but with modified material parameters. Introducing $\tilde{\epsilon}=\epsilon[1+i\mathrm{Im}(\omega)/(\omega)]$ and $\tilde{\mu}=\mu_{0}[1+i\mathrm{Im}(\omega)/(\omega)]$, we can show that [34]
@@
where the fields should now be understood to be oscillating at the real frequency $(\omega)$. Note that the non-zero imaginary parts of $\tilde{\epsilon}$ and $\tilde{\mu}$ introduce virtual gain or loss to the system, which is an artifact of the fact that $\omega$ is actually complex. Eq. (6) is an energy balance relation for the system, since the integral on the right hand side describes the energy dissipated (gained) within the system by the virtual loss (gain) [39]. $B_{\omega}$ meanwhile is a complex quantity whose real part is given by [34]
@@
which, up to constant factors, is equal to the system’s stored electromagnetic energy [39]. Frequency derivatives in Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation) are necessary to properly account for the effects of material dispersion. On the other hand, $\mathrm{Im}(B_{\omega})=-\partial_{(\omega)}A/2$ describes energy dissipation within the system. Similar expressions have been used to account for material losses in resonant systems [40, 41].
Consider now $B_{\alpha}$ and suppose that the system perturbation has the effect of modifying the permittivity in a localized region $\Omega_{\alpha}\subset\Omega$. This might be caused by, for example, a local pressure change or a particle binding to the system. If the system’s internal field distributions are only weakly affected by the perturbation, then $V^{e}_{\alpha}=V^{m}_{\alpha}\approx 0$ and we have
@@
Recalling the form of the right hand side of Eq. (4), note that the numerator will contain the factor $\Delta\alpha B_{\alpha}$, which, in light of Eq. (8), will involve the product $\Delta\alpha\partial_{\alpha}\epsilon$. By assumption, outside of $\Omega_{\alpha}$, $\partial_{\alpha}\epsilon=0$, while within $\Omega_{\alpha}$ we have $\Delta\alpha\partial_{\alpha}\epsilon=\Delta\epsilon$, where $\Delta\epsilon$ is the change in $\epsilon$ caused by the perturbation. Eq. (4) therefore reduces to
@@
Since we have assumed high quality factor resonances, the previously discussed virtual gain or dissipation will be weak, such that $B_{\omega}\approx(B_{\omega})$ and Eq. (9) is thus in full agreement with standard cavity perturbation theory [24]. It is important to stress that Eq. (9) is only strictly valid for infinite quality factor resonances, but is a reasonable approximation when loss is weak. In contrast, Eq. (3) holds more generally, since no restrictive assumptions were made in its derivation. For further analysis of the case of low quality factor resonances, we refer the interested reader to Ref. [26].
Numerical examples.—We now turn to demonstrating the validity of our results with numerical simulations. As an example system, we consider photonic networks consisting of randomly connected, single-mode dielectric waveguides, which have recently been investigated as a platform for random lasing [37] and in integrated photonic circuits [42]. Such systems are relatively simple to model and have non-trivial spectra. An example network, shown in Figure 1, was generated by Delaunay triangulation of a random collection of coplanar points (the purple, or ‘internal’ nodes). An additional layer of ‘external’ nodes (green) connect to some of the internal nodes at the outer edges of the network and serve as entry points to the network, allowing one to define the network scattering matrix $\mathbf{S}$. Given knowledge of the scattering properties of the network’s nodes and links, $\mathbf{S}$ and its derivatives can be calculated and thus $\mathbf{Q}_{\xi}$ evaluated directly [43]. In the supplemental material [34], we also demonstrate how $\mathbf{Q}_{\xi}$ can be calculated numerically from Eq. (5).
@@
With our network, we searched for scattering matrix poles with optical frequencies $(\omega)$ in the vicinity of $3425\,\mathrm{THz}$, corresponding to a vacuum wavelength of about $550\,\mathrm{nm}$. The links were assumed to be made of BK7 glass with refractive index $n$ calculated using a standard Sellmeier equation [44]. Dispersion was weak in our example and $n\approx 1.5185$ was approximately constant over the range of frequencies considered. Propagation of light through each link was described by exponential factors of the form $e^{\pm i\omega nL/c}$, where $L$ is the link length. Our network was micro-scale in size with an average link length of $60\,\mu$m. For simplicity, each internal node was given a randomly generated, frequency-independent scattering matrix drawn from the circular orthogonal ensemble to enforce energy conservation and reciprocity [45]. Although not fully realistic, our model was sufficient to produce a complex scattering system with which our theory could be tested. To perturb the network, we isolated a segment of a randomly chosen link (the red, ‘perturbed’ segment or $\Omega_{\alpha}$ in Figure 1) and varied its refractive index relative to the rest of the link. Specifically, the refractive index of the segment $n_{s}$ was given by $n_{s}=n+\Delta n$, where $\Delta n$ was incrementally increased from $10^{-5}$ to $10^{-2}$. Note that the final value $\Delta n=10^{-2}$ corresponded to a total phase shift in the segment $\phi=\omega\Delta nL/c$ equal to several multiples of $2\pi$. We further introduced two virtual nodes where the perturbed segment met the rest of the link, which were given standard Fresnel scattering matrices based on the refractive index mismatch.

### v2

@@ Abstract
Introduction—Resonance is a fundamental physical phenomenon that plays a key role in fields as diverse as quantum mechanics [1], structural engineering [2], electromagnetics [3], and fluid dynamics [4]. In closed systems, resonances are associated with normal modes, which correspond to the orthogonal eigenvectors of the system’s Hermitian Hamiltonian. The corresponding real eigenvalues define the spectrum of permissible oscillation frequencies or energy levels [5]. Eigenvalue problems, however, generally do not admit analytical solutions. Approximate methods, in which a system is modelled by perturbing a solvable reference problem, are therefore frequently employed. Rayleigh-Schrödinger perturbation theory, in which the perturbed solution is written as a series expansion over all unperturbed eigenstates, is perhaps the most famous approach [6], however a multitude of alternate theories have been developed, such as Brillouin-Wigner, many-body or lattice perturbation theory [7, 8, 9].
@@
Although closed systems are useful theoretical ideals, in reality a system interacts with its environment. The coupling of normal modes to external degrees of freedom introduces potential loss channels, which can alter the nature and energy of resonant modes. So-called open, or scattering, systems can no longer be described by a Hermitian Hamiltonian and a coupling operator must be introduced to form an effective non-Hermitian Hamiltonian [10]. Alternatively, one can use the closely related scattering matrix, $\mathbf{S}$, which derives from the resolvent of the effective Hamiltonian [11]. The resonant states of the system, now termed quasi-normal modes (QNMs), are then identified from the poles of $\mathbf{S}$ when analytically continued into the complex frequency plane [12, 5]. Notably, loss pushes the resonant frequencies off the real axis, resulting in complex-valued eigenfrequencies $\omega_{p}$, where $(\omega_{p})$ and $-\mathrm{Im}(\omega_{p})/2$ are the resonant frequency and linewidth.
@@
Open systems are known to possess a number of interesting resonant and scattering phenomena, such as coherent perfect absorption [13], exceptional points [14], and bound states in the continuum [15]. Moreover, resonances in open systems are susceptible to environmental variations, such as changes in temperature, external fields or material composition, a concept underpinning modern sensors [16, 17, 18]. Perturbed non-Hermitian systems can exhibit additional unique behavior, such as anomalous resonance shifts [19, 20] and nonlinear sensitivity [21]. Evaluation of resonant modes and their properties therefore remains an important task in, for example, evaluating state lifetimes, laser dynamics and sensor sensitivity. Traditional perturbation theories, however, are not directly applicable to open systems due to mathematical complexities introduced by QNMs [22, 23]. For high quality factor resonances, such as those supported by optical microcavities [24], coupling to the environment is weak allowing resonance shifts to be expressed as changes in stored energy [25, 26]. The validity of such expressions is however limited and they fail to correctly predict changes in linewidths [27], which is now understood to be a signature of the non-orthogonality of QNMs [28, 29]. In the presence of significant losses, more careful consideration of the normalization of QNMs [27] or alternative mode expansions [30] become necessary.
In this paper, we present a novel formulation for determining resonance shifts due to arbitrary system perturbations based on generalized Wigner-Smith (WS) operators. In the case of high quality factor resonances, we show that these operators are related to a system’s stored and dissipated energy, in full agreement with results from traditional perturbation theory. Our theory is tested against numerical simulations of resonant scattering in random photonic networks. In an accompanying paper [31], we apply our theory to low quality factor resonances and present further numerical examples.
@@
Perturbation theory—An important tool commonly employed in the analysis of scattering systems is the WS time delay matrix
@@
where $\partial_{\omega}$ denotes the partial derivative with respect to frequency $\omega$ [32, 33]. $\mathbf{Q}_{\omega}$ has several interesting properties relevant to resonant scattering. For unitary systems at real $\omega$, $\mathbf{Q}_{\omega}$ is Hermitian and its eigenvalues are associated with well-defined time delays experienced by narrow-band, transient system excitations. Evaluated at scattering matrix poles, these time delays coincide with the lifetimes of the associated resonances, demonstrating a link between $\mathbf{Q}_{\omega}$ and a system’s QNMs [34]. In electromagnetic theory, it has also been shown that diagonal elements of $\mathbf{Q}_{\omega}$ can be expressed as energy-like integrals over the extent of the system [35], highlighting the connection to mode volume.
By itself $\mathbf{Q}_{\omega}$ lacks the specificity required to efficiently capture localized or parametric system perturbations. For this reason, we also consider the generalized WS operator $\mathbf{Q}_{\alpha}$, defined by Eq. (1) with $\omega\rightarrow\alpha$, where $\alpha$ is an arbitrary variable. Operators of this kind have recently attracted attention in the optical domain as a tool for engineering light in complex scattering environments [36, 37, 38]. Importantly, by virtue of its generalized nature, $\mathbf{Q}_{\alpha}$ contains sufficient information to predict resonance shifts induced by arbitrary system perturbations. To show this, we consider a generic scattering system described by a scattering matrix $\mathbf{S}$, whose poles are assumed to be of order one. Let $\alpha$ ($\alpha^{\prime}$) denote some system parameter, such as the temperature, shape or size of some component of the system, before (after) it is perturbed. To determine the change in the resonant frequency $\Delta\omega_{p}=\omega_{p}^{\prime}-\omega_{p}$ that is induced by the perturbation $\Delta\alpha=\alpha^{\prime}-\alpha$, consider the function $f(\omega,\alpha)=\det[\mathbf{S}^{-1}(\omega,\alpha)]$, whose zeros coincide with the poles of $\mathbf{S}$. Assuming the perturbation is small, a first order, multivariate Taylor expansion of $f$ gives
@@
where tr denotes the trace operator and $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}$ denotes the residue at the pole $\omega_{p}$. As can be seen from Eq. (3), for a given perturbation $\Delta\alpha$, the pole shift $\Delta\omega_{p}$ is completely determined by $\mathop{\mathrm{Res}}_{\omega=\omega_{p}}\mbox{tr}(\mathbf{Q}_{\alpha})$, which should be evaluated for the unperturbed network. We also observe that the form of Eq. (3) implies the ratio of the residues of the generalized WS operators behaves as a condition number for $\omega_{p}$ when treated as a function of $\alpha$ [40]. Eq. (3) is our key result, which, importantly, was derived on purely mathematical grounds, and is therefore applicable to a broad range of physical scenarios.
@@
To gain physical insight into Eq. (3), we consider now a class of optical systems composed of arbitrary, finite interior regions coupled to the environment by a finite number of dielectric waveguides. Examples of such systems include fiber Bragg resonators, complex networks and ring resonators [41, 42, 43]. Note that Eq. (3) can be written in the form
@@
$\Delta\omega_{p}=\lim_{\omega\to\omega_{p}}\left[-\Delta\alpha\frac{\mbox{tr}(\mathbf{Q}_{\alpha})}{\mbox{tr}(\mathbf{Q}_{\omega})}\right],$ (4)
@@
and so the pole shift is given approximately by the expression within the limit evaluated at $\omega\approx\omega_{p}$. We assume for simplicity that the system permittivity $\epsilon$ is real and isotropic and the permeability $\mu_{0}$ is that of free space. Restricting to resonances have high quality factors, such that $\mathrm{Im}(\omega)\ll(\omega)$, it is possible to show that [39]
$\mathbf{Q}_{\xi}$ $=\left[\mathbf{I}+2\mathrm{Im}(\omega)\int_{\Omega}(\epsilon\mathbf{U}^{e}+\mu_{0}\mathbf{U}^{m})\,\mathrm{d}V\right]^{-1}$
$\quad\quad\bigg{[}-\int_{\Omega}\big{[}\partial_{\xi}(\omega\epsilon)\mathbf{U}^{e}+\partial_{\xi}\omega\mu_{0}\mathbf{U}^{m}$
$\quad\quad\quad\quad\quad+2i\mathrm{Im}(\omega)(\epsilon\mathbf{V}_{\xi}^{e}+\mu_{0}\mathbf{V}_{\xi}^{m})\big{]}\,\mathrm{d}V\bigg{]},$ (5)
@@
where $\xi$ represents either $\omega$ or $\alpha$ and $\Omega$ denotes the volume occupied by the system with a boundary $\partial\Omega$ perforated only by the coupling waveguides. In Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation), $\mathbf{I}$ is the identity matrix and $\mathbf{U}^{e},\mathbf{U}^{m},\mathbf{V}^{e}_{\xi}$, and $\mathbf{V}^{m}_{\xi}$ are matrices whose $(q,p)$–th elements are given by $U^{e}_{qp}=\frac{1}{4}\mathbf{E}_{p}\cdot\mathbf{E}^{*}_{q}$, $U^{m}_{qp}=\frac{1}{4}\mathbf{H}_{p}\cdot\mathbf{H}^{*}_{q}$, $V^{e}_{\xi,qp}=\frac{1}{4}\partial_{\xi}\mathbf{E}_{p}\cdot\mathbf{E}^{*}_{q}$, and $V^{m}_{\xi,qp}=\frac{1}{4}\partial_{\xi}\mathbf{H}_{p}\cdot\mathbf{H}^{*}_{q}$ respectively, where the fields $\mathbf{E}_{p}$ and $\mathbf{H}_{p}$ ($\mathbf{E}_{q}$ and $\mathbf{H}_{q}$) are those that exist throughout $\Omega$ when the system is illuminated by the $p$–th ($q$–th) incident field. Here, $p$ and $q$ enumerate all modes in all waveguides that couple the system to the environment.
@@
As discussed in the Supplemental Material [39], the form of $\mathbf{Q}_{\xi}$ in Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation) is linked to the factorization $\mathbf{Q}_{\xi}=\mathbf{A}^{-1}\mathbf{B}_{\xi}$, where $\mathbf{A}=\mathbf{S}^{\dagger}\mathbf{S}$ and $\mathbf{B}_{\xi}=-i\mathbf{S}^{\dagger}\partial_{\xi}\mathbf{S}$. We shall analyze these factors separately. First, to further simplify matters, we consider the case where the system has a single coupling waveguide supporting a single mode, so that the scattering matrix reduces to an effective scalar reflection coefficient $r_{\mathrm{eff}}$. Correspondingly, the WS operator $\mathbf{Q}_{\xi}$ reduces to the scalar quantity $Q_{\xi}=A^{-1}B_{\xi}$, where $A=|r_{\mathrm{eff}}|^{2}$ and $B_{\xi}=-ir^{*}_{\mathrm{eff}}\partial_{\xi}r_{\mathrm{eff}}$. To evaluate these expressions at complex $\omega$, we recast the problem into one at a real frequency, but with modified material parameters. Introducing $\tilde{\epsilon}=\epsilon[1+i\mathrm{Im}(\omega)/(\omega)]$ and $\tilde{\mu}=\mu_{0}[1+i\mathrm{Im}(\omega)/(\omega)]$, we can show that [39]
@@
where the fields should now be understood to oscillate with real frequency $(\omega)$. Note, the non-zero imaginary parts of $\tilde{\epsilon}$ and $\tilde{\mu}$ introduce virtual gain or loss to the system, which is an artifact of the fact that $\omega$ is actually complex. Eq. (6) is an energy balance relation, since the integral on the right hand side describes the energy dissipated (gained) within the system by the virtual loss (gain) [44]. $B_{\omega}$ meanwhile is complex with [39]
@@
which, up to constant factors, equals the system’s stored electromagnetic energy [44]. Frequency derivatives in Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation) are necessary to account for material dispersion. On the other hand, $\mathrm{Im}(B_{\omega})=-\partial_{(\omega)}A/2$ describes energy dissipation within the system. Similar expressions have been used to account for material losses in resonant systems [45, 46].
Consider now $B_{\alpha}$ and suppose that the system perturbation modifies the permittivity in a localized region $\Omega_{\alpha}\subset\Omega$. This might be caused by, for example, a local pressure change or a particle binding to the system. If the internal field distributions are only weakly affected by the perturbation, then $V^{e}_{\alpha}=V^{m}_{\alpha}\approx 0$ and
@@
Recalling the form of the right hand side of Eq. (4), note that the numerator will contain the factor $\Delta\alpha B_{\alpha}$, which, in light of Eq. (8), will involve the product $\Delta\alpha\partial_{\alpha}\epsilon$. By assumption, outside of $\Omega_{\alpha}$, $\partial_{\alpha}\epsilon=0$, while within $\Omega_{\alpha}$ we have $\Delta\alpha\partial_{\alpha}\epsilon=\Delta\epsilon$, where $\Delta\epsilon$ is the change in $\epsilon$. Eq. (4) therefore reduces to
@@
Since we have assumed high quality factor resonances, the virtual gain or dissipation will be weak, such that $B_{\omega}\approx(B_{\omega})$ and Eq. (9) is thus in full agreement with standard cavity perturbation theory [27]. It is important to stress that Eq. (9) is only strictly valid for infinite quality factor resonances, but is a reasonable approximation when loss is weak. In contrast, Eq. (3) holds more generally, since no restrictions were made in its derivation. Further analysis of low quality factor resonances can be found in Ref. [31].
Numerical examples—We now turn to demonstrating the validity of our results with numerical simulations. As an example, we consider photonic networks consisting of randomly connected, single-mode dielectric waveguides, which have recently been investigated as a platform for random lasing [42, 47] and in integrated photonic circuits [48]. Such systems are relatively simple to model and have non-trivial spectra. An example network, shown in Figure 1, was generated by Delaunay triangulation of a random collection of coplanar ‘internal’ nodes (purple). An additional layer of ‘external’ nodes (green) connect to selected internal nodes at the network periphery and serve as entry points, allowing one to define the network scattering matrix $\mathbf{S}$. Given knowledge of the scattering properties of the network’s nodes and links, $\mathbf{S}$ and its derivatives can be calculated and thus $\mathbf{Q}_{\xi}$ evaluated directly [49]. In the Supplemental Material [39], we demonstrate how $\mathbf{Q}_{\xi}$ can also be found numerically from Eq. (Perturbing scattering resonances in non-Hermitian systems: a generalized Wigner-Smith operator formulation).
@@
With our network, we searched for poles of $\mathbf{S}$ with frequencies $(\omega)\sim 3425\,\mathrm{THz}$, corresponding to a vacuum wavelength of $\sim 550\,\mathrm{nm}$. The links were assumed to be made of glass with refractive index $n$ calculated using a standard Sellmeier equation [50]. Dispersion was weak in our example and $n\approx 1.5185$ was approximately constant over the frequency range considered. Propagation of light through each link was described using exponential factors of the form $e^{\pm i\omega nL/c}$, where $L$ is the link length. The average link length in the network was $60\,\mu$m. For simplicity, each internal node was given a randomly generated, frequency-independent scattering matrix drawn from the circular orthogonal ensemble to enforce energy conservation and reciprocity [51]. To perturb the network, we isolated a segment of a randomly chosen link (red ‘perturbed’ segment, $\Omega_{\alpha}$, in Figure 1) and varied its refractive index $n_{s}$. Specifically, $n_{s}$ was given by $n_{s}=n+\Delta n$, where $\Delta n$ was incrementally increased from $10^{-5}$ to $10^{-2}$. Note that the final value $\Delta n=10^{-2}$ corresponded to a total phase shift in the segment $\phi=\omega\Delta nL/c$ equal to several multiples of $2\pi$. We further introduced two virtual nodes at the ends of the perturbed segment, which were given standard Fresnel scattering matrices based on the refractive index mismatch.